OpenStudy (anonymous):

write \(\large (1-i)^{2i}\) as of the form x+iy

OpenStudy (anonymous):

@ganeshie8 wanna try ?

OpenStudy (anonymous):

@Loser66

ganeshie8 (ganeshie8):

\[\large (1-i)^{2i} = e^{2i \ln (1-i)}\]

OpenStudy (anonymous):

how ?and sure :O when there is i im not sure of log or ln

ganeshie8 (ganeshie8):

lets hope it works... :) we can check wid wolfram in the end maybe

OpenStudy (anonymous):

log z = ln r +i ( theta +2k pi)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

i never took complex analysis so idk from where u pulled up that formula

ganeshie8 (ganeshie8):

oh ok got it :)

OpenStudy (anonymous):

from the notes there is box full of formulas xD

ganeshie8 (ganeshie8):

okay looks like the log thingy works : http://www.wolframalpha.com/input/?i=%281-i%29%5E%282i%29++%3D+e%5E%282i*ln%281-i%29%29

OpenStudy (anonymous):

ok find somthing log z = ln |z| + i arg z

OpenStudy (anonymous):

ok go ahead im doing this with music https://www.youtube.com/watch?v=fSKCJrh_VRc

ganeshie8 (ganeshie8):

\[\large (1-i)^{2i} = e^{2i \ln (1-i)} = e^{2i \ln \left(\sqrt{2}e^{i(\frac{-\pi}{4})}\right)}\]

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

no , log e^z \(\neq\) z :\

OpenStudy (anonymous):

but e^log z = z

ganeshie8 (ganeshie8):

why not

OpenStudy (anonymous):

one of the properties

OpenStudy (anonymous):

log e^z =z + i2k pi

ganeshie8 (ganeshie8):

well we're using the second property, so we should be fine ?

OpenStudy (anonymous):

ok keep going :O but i dont trust walfram sometimes

ganeshie8 (ganeshie8):

\[\large (1-i)^{2i} = e^{2i \ln (1-i)} = e^{2i \ln \left(\sqrt{2}e^{i(\frac{-\pi}{4})}\right)} = e^{2i \left[\ln \sqrt{2} - \frac{i\pi}{4}\right] }\]

OpenStudy (anonymous):

nw we have to convert to x+iy

ganeshie8 (ganeshie8):

yes break them

OpenStudy (anonymous):

r=2 theta=the rest ?

ganeshie8 (ganeshie8):

break it into two using product property of exponents : \(a^{m+n} = a^m.a^n\)

OpenStudy (anonymous):

O.O

ganeshie8 (ganeshie8):

\[\large (1-i)^{2i} = e^{2i \ln (1-i)} = e^{2i \ln \left(\sqrt{2}e^{i(\frac{-\pi}{4})}\right)} = e^{2i \left[\ln \sqrt{2} - \frac{i\pi}{4}\right] } \] \[\large = e^{2i \ln \sqrt{2} } . e^ {2i\frac{-i\pi}{4}} = e^{i \ln 2 } . e^ {\frac{\pi}{2}} \]

OpenStudy (anonymous):

ok then ?

ganeshie8 (ganeshie8):

we're done

ganeshie8 (ganeshie8):

\[\large (1-i)^{2i} = e^{2i \ln (1-i)} = e^{2i \ln \left(\sqrt{2}e^{i(\frac{-\pi}{4})}\right)} = e^{2i \left[\ln \sqrt{2} - \frac{i\pi}{4}\right] } \] \[\large = e^{2i \ln \sqrt{2} } . e^ {2i\frac{-i\pi}{4}} = e^{i \ln 2 } . e^ {\frac{\pi}{2}} \] \[ = e^ {\frac{\pi}{2}} \left[ \cos(\ln 2 ) + i\sin (\ln 2) \right] \]

OpenStudy (anonymous):

ok lol i was confused , but nw only i got it xD ty !

OpenStudy (anonymous):